Data for the next 2 questions: Kim Companys utility bill is $800 per month when there is no production (fixed utility cost
Answer #1

5.

Fixed utilities cost per month = $800

Utility cost per unit = $1

Total utility cost = (Fixed utilities cost per month + Utility cost per unit) x Number of units

1,000 units 1,500 units 2,000 units
Total utility cost 800 + 1,000 x 1 = $1,800 800 + 1,500 x 1 = $2,300 800 + 2,000 x 1 = $2,800

6.

1,000 units 1,500 units 2,000 units
Utility cost per unit (Total utility cost/Number of units) 1,800/1,000 = $1.8 2,300/1,500 = $1.53 2,800/2,000 = $1.4

7.

Fixed manufacturing overhead per unit when 4,000 bicycle are produced = $12.50

Total fixed manufacturing overhead at 4,000 bicycle = 4,000 x 12.50

= $50,000

Total fixed manufacturing overhead will not change due to change in output. Hence, when output is increased to 5,000 bicycle, total fixed manufacturing overhead will remain $50,000.

Variable cost per unit will remain same even due to increase in output.

Direct material 25
Direct manufacturing labor 5
Variable manufacturing overhead 14.50
Fixed manufacturing overhead (50,000.5,000) 10
Total manufacturing cost $54.50

kindly give a positive rating if you are satisfied with the solution. Please ask if you have any query related to the question, Thanks.